# Robust and Fine-Grained Detection of AI Generated Texts

### v2

@@ 6 Other Observations
The sentences inside which text authorship switches from human to LLM or vice versa were found to be relatively shorter that the original text portions which they replaced. LLMs may be likely to finish the current sentence earlier than usual to move on to the next sentence in text completion scenarios. The mean length of the original portion and the replaced portions of those sentences for each language and generator can be seen in Table 5 and Table 6 respectively. This observation was consistent across all languages and generators with a 20-30% reduction and a larger reduction in Hindi. For Chinese and Japanese too, we did observe a 20-30% reduction in character count when comparing the original and replaced portions of the sentence after the text boundary. Although there is a good variation in this feature across languages, the mean and medians observed for each language were similar for all the LLMs. This is further elaborated in Appendix A.
@@ 7 Conclusion
Despite not being trained on the domains or generators, the models built through our approach performed well over several benchmarks as seen in subsection 5.3 and subsection 5.4 over inputs which were from non-native speakers and consist of adversarial methods. Further, one case where many proprietary systems struggle is when the inputs were too short, which our models were able to overcome as seen in Figure 6, which demonstrates our models’ accuracy over our test set compared to input text’s sentence count. Table 7 displays our model’s performance over English subset of our dataset for each generator. A similar trend from subsection 5.4 was observed with models which are likely less instruction-tuned / not instruction-tuned tend to produce texts which are harder to distinguish than their alternatives.
@@ 7.1 Scalability and scope for extension
The original dataset used to train our current models as mentioned in section 3 consists of samples over 60 languages which would cover 70% of the world population’s primary language, and all of the languages are supported by existing multilingual transformer models making the process of scaling the work to more languages easier. Despite not being trained on the generators or domains’ texts, our models were able to perform well on several benchmarks. Even reaching a F1 score of 0.79 against adversarial inputs while they were neither trained over them nor pre-processed. Similarly, creation and usage of such large datasets of other domains along with ours might result in robust and better models. We couldn’t explore the relation between instruction tuning sample size of LLMs and detectability of their texts due to the proprietary nature of most of the generators we used, but a similar study using open-data models could uncover more insights.
@@ 7.2 Scope for Improvement
As seen in Figure 5, almost none of the adversarial methods affected the models built through our approach other than paraphrasing and homo-glyphs. However homo-glyphs can be pre-processed by mapping them to the actual character they were imitating in the text. This would require a large collection of homo-glyph to character mapping set to use for pre-processing. Further, paraphrased samples of various number of iterations being included in the training dataset might lead to further improvements. It is also worth exploring how detectable are texts in cases where multiple generators contribute a portion each in a human authored text. Other missing adversarial methods that are likely to be used in practical scenarios include usage of proprietary systems that ’humanize’ a given text in an attempt to evade detection.
7.3 Ideal Usage
The models were built primarily for a human-in-the-loop use cases where the model would try to flag most of the likely machine-generated portions while the flagged content can be validated either through an ensemble of models or a human and hence a tilt towards higher recall can be observed in the metrics as seen in Table 13.
@@ 7.3 Ideal Usage
Just like any other detector or classifier, no detector can guarantee a 100% accuracy and hence the models are not meant to be used directly for decision making but are meant to be used in a human-in-the-loop scenarios. Furthermore, the experiments carried out did not include cases of multiple LLMs co-authoring a portion each of the same text.
Acknowledgements
Several stages of this work were supported by grants from Traversaal.ai, Cohere Labs & Microsoft Azure.
@@ Appendix C Reproducibility
We used multilingual longformer  [^7] https://huggingface.co/hyperonym/xlm-roberta-longformer-base-16384 with an additional CRF layer. The hyper-parameters used for training the models can be seen in Table 10. We built a separate model for each language, the training was done over H100 SXM over 10h each.
@@ Appendix D Other Metrics
Appendix E License
The xlm-longformer model we used was available with an mit license, we are releasing the models and datasets through CC BY-NC 4.0888https://creativecommons.org/licenses/by-nc/4.0/deed.en which permits usage for research purposes.

### v3

@@ 6 Other Observations
The sentences inside which text authorship switches from human to LLM or vice versa were found to be relatively shorter than the original text portions that they replaced. LLMs may be likely to finish the current sentence earlier than usual to move on to the next sentence in text completion scenarios. The mean length of the original portion and the replaced portions of those sentences for each language and generator can be seen in Table 6 and Table 7, respectively. This observation was consistent across all languages and generators with a 20-30% reduction and a larger reduction in Hindi. For Chinese and Japanese too, we did observe a 20-30% reduction in character count when comparing the original and replaced portions of the sentence after the text boundary. Although there is a good variation in this feature across languages, the mean and medians observed for each language were similar for all the LLMs. This is further elaborated in Appendix A.
@@ 7 Conclusion
Despite not being trained on the domains or generators, the models built through our approach performed well over several benchmarks as seen in subsection 5.3 and subsection 5.4 over inputs that were from non-native speakers and consist of adversarial methods. Further, one case where many proprietary systems struggle is when the inputs were too short, which our models were able to overcome as seen in Figure 6, which demonstrates our models’ accuracy over our test set compared to the input text’s sentence count. Table 8 displays our model’s performance over the English subset of our dataset for each generator. A similar trend from subsection 5.4 was observed with models that are likely less instruction-tuned or not instruction-tuned; they tend to produce texts that are harder to distinguish than their alternatives.
@@ 7.1 Scalability and scope for extension
The original dataset used to train our current models, as mentioned in section 3, consists of samples in over 60 languages, which would cover 70% of the world population’s primary language, and all of the languages are supported by existing multilingual transformer models, making the process of scaling the work to more languages easier. Despite not being trained on the generators or domains’ texts, our models were able to perform well on several benchmarks. They even reached an F1 score of 0.79 against adversarial inputs while they were neither trained over them nor pre-processed. Similarly, the creation and usage of such large datasets of other domains along with ours might result in robust and better models. We couldn’t explore the relation between the instruction tuning sample size of LLMs and the detectability of their texts due to the proprietary nature of most of the generators we used, but a similar study using open-data models could uncover more insights.
@@ 7.2 Scope for Improvement
As seen in Figure 5, almost none of the adversarial methods affected the models built through our approach other than paraphrasing and homoglyphs. However, homoglyphs can be pre-processed by mapping them to the actual character they were imitating in the text. This would require a large collection of homo-glyph-to-character mapping sets to use for preprocessing. Further, paraphrased samples of various numbers of iterations being included in the training dataset might lead to further improvements. It is also worth exploring how detectable texts are in cases where multiple generators contribute a portion each in a human-authored text. Other missing adversarial methods that are likely to be used in practical scenarios include the usage of proprietary systems that ’humanize’ a given text in an attempt to evade detection.
@@
Just like any other detector or classifier, no detector can guarantee a 100% accuracy and hence the models are not meant to be used directly for decision making but are meant to be used in a human-in-the-loop scenarios. Furthermore, the experiments carried out did not include cases of multiple LLMs co-authoring a portion each of the same text. The models were built primarily for a human-in-the-loop use cases where the model would try to flag most of the likely machine-generated portions while the flagged content can be validated either through an ensemble of models or a human and hence a tilt towards higher recall can be observed in the metrics
@@ Appendix C Reproducibility
We used multilingual longformer  [^8] https://huggingface.co/hyperonym/xlm-roberta-longformer-base-16384 with an additional CRF layer. The hyper-parameters used for training the models can be seen in Table 11. We built a separate model for each language, the training was done over H100 SXM over 237 hours.